OpenStudy (anonymous):

Find the volume of a cylinder with a diameter of 8 inches and a height that is three times the radius. Use 3.14 for pi and round your answer to the nearest hundredth. (Hint: You may only enter numerals, decimal points, and negative signs in the answer blank)

OpenStudy (sleepyhead314):

V = (1/3)pi r^2 h r = 8/2 = 4 h = 3 r = 3*4 = 12

OpenStudy (anonymous):

602.88?

OpenStudy (king.void.):

http://openstudy.com/updates/5283cc3ee4b0218bc6510869

OpenStudy (anonymous):

Oh wow void XD isn't that kinda cheating? :P

OpenStudy (sleepyhead314):

oh whoops xD gave the wrong equation but yes, your answer is correct @Tatsurion :) good job :D <3

OpenStudy (king.void.):

It's on OpenStudy, so no :P and he explains it

OpenStudy (anonymous):

<3 :DDDDD

OpenStudy (sleepyhead314):

and @King.Void. it is actually considered cheating... "any explanation which does not involve the askers participation, but does all the work/thinking for them ... is nothing more than a direct answer. if doesnt matter if your explanation is 3 pages long and the most indepth textual treatise on the planet; without the askers participation all they do is say ok, take the final results, and move on to ask the same exact question all over again." source: http://openstudy.com/users/igreen#/updates/5444f6ffe4b044d2f3c2a2a2

OpenStudy (king.void.):

sorry :/